OpenStudy (anonymous):

You are looking down at a map. A vector U with ||U||= 7 points north and a vector V with ||V|| = 8 points northeast. The cross product U X V points ; What is the magnitude ||U X V||=??

OpenStudy (anonymous):

CMIIW

OpenStudy (anonymous):

\[|| U \times V || = \sqrt{7^{2}+8^{2}-2(7)(8)}\]

OpenStudy (anonymous):

i got 1 and its not correct

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

maybe my formula was wrong whats your formula

OpenStudy (anonymous):

i had similar to your just not the -2(7)(8)

OpenStudy (anonymous):

please the equation writes

OpenStudy (anonymous):

\[|| U \times V || = \sqrt{U ^{2}+V ^{2}-2UV}\]

OpenStudy (anonymous):

like this ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

hello